OpenStudy (anonymous):

A car rent for $35 per day plus 21cents per mile. You are on a daily budget of $98. What mileage will allow you to stay within your budget?

OpenStudy (anonymous):

miles allowed=(98-35)/0.21=63/0.21=300miles